OpenStudy (ambermarie151):

What is the solution? Y=-2x-4 Y=x+2 A.(0,2) B.(0,-4) C(-2,-4) D.(-2,0)

OpenStudy (aussiebabe23):

Set the equations equal to each other and solve for x

OpenStudy (mathmate):

or, rewrite the first equation as y=-2(x+2) and compare with y=x+2 and conclude the solution for y, hence your choice.

OpenStudy (aussiebabe23):

-2x-4=x+2 -2x-6=x -6=3x x=-2

OpenStudy (ambermarie151):

So it would be C

OpenStudy (aussiebabe23):

plug -2 in for x for one of the equations and see what you get, that will be the y value

OpenStudy (ambermarie151):

I'm confused

OpenStudy (aussiebabe23):

its D

OpenStudy (welshfella):

both -2x - 4 and x + 2 are equal to Y therefore the 2 expressions ion x are equal to each other so we can write -2x - 4 = x+ 2 solve for x

OpenStudy (welshfella):

gtg

OpenStudy (mathmate):

@aussiebabe23 I would have loved to see you explain y=x+2=(-2)+2=0, so the solution is (x,y)=(-2,0). @ambermarie151 can then at least make her choice of the answer. The fact that she's already offline means she only wanted the answer, not how to get it. Sigh.

OpenStudy (aussiebabe23):

i did explain... did you not read the post prior to that. If the girl doesn't want to learn you can't force her to read. let her get a high school diploma so she can at least get a job as a janitor

OpenStudy (mathmate):

@aussibabe23 I agree you did try, and I appreciate it. However, I have seen so much of this "get answer and run" that I do not give out the letter answers. Even if I work out the correct answer, I do not provide the direct answer choice (in letters). This way, the asker will at least have to take the responsibility of make the right choice. Thank you for the good work, and it seems that you understand the learning problem of today.

OpenStudy (mathmate):

@aussiebabe23 sorry, got your name wroing. See previous post.